# Vision-based Manipulation from Single Human Video with Open-World Object Graphs

### v1

@@ Abstract
We present an object-centric approach to empower robots to learn vision-based manipulation skills from human videos. We investigate the problem of imitating robot manipulation from a single human video in the open-world setting, where a robot must learn to manipulate novel objects from one video demonstration. We introduce ORION, an algorithm that tackles the problem by extracting an object-centric manipulation plan from a single RGB-D video and deriving a policy that conditions on the extracted plan. Our method enables the robot to learn from videos captured by daily mobile devices such as an iPad and generalize the policies to deployment environments with varying visual backgrounds, camera angles, spatial layouts, and novel object instances. We systematically evaluate our method on both short-horizon and long-horizon tasks, demonstrating the efficacy of ORION in learning from a single human video in the open world. Videos can be found in the project website.
@@ 1 Introduction
A critical step toward building robot autonomy is developing sensorimotor skills for perceiving and interacting with unstructured environments. Conventional methods for acquiring skills necessitate manual engineering and/or costly data collection [1, 2, 3, 4, 5]. A promising alternative is teaching robots through human videos of manipulation behaviors situated in everyday scenarios. These methods have great potential to tap into the readily available source of Internet videos that encompass a wide distribution of human activities, paving the ground for scaling up skill learning.
@@
Prior work on learning from human videos has focused on pre-training representations and value functions [6, 7, 8, 9, 10]. However, they do not explicitly capture object states and their interactions in 3D space where robot motions are defined. Consequently, they require separate teleoperation data for each set of objects in each location and even for each possible change in visual background, e.g., the scene background or lighting conditions [11]. In contrast, our goal is for a robot to learn to perform a task robustly in the “open world”, i.e., under varying visual and spatial conditions from a single human video, without prior knowledge of the object models or the behaviors shown. Since our policy construction process uses actionless videos that are equivalent to state-only demonstrations in the problem of “Imitation from Observation”[12], we refer to our problem setting as open-world imitation from observation.
@@
Developing a method in this setting is only possible due to the recent advances in vision foundation models [13, 14]. These models, pre-trained on Internet-scale visual data, excel at understanding open-vocabulary visual concepts and enable robots to recognize and localize objects in natural videos without known object categories or access to physical states. This work marks the first step toward achieving our vision of open-world imitation from observation, where a robot learns to interact with objects given a single video while generalizing to environments with different visual backgrounds and unseen spatial configurations during deployment. In this work, we consider using RGB-D video demonstrations where a person manipulates a small set of task-relevant objects with their single hand, recorded with a stationary camera. These videos are actionless or state-only, as they do not come with any ground-truth action labels for the robot.
@@
Figure 1: Overview. We introduce ORION for tackling the problem of learning manipulation behaviors from single human video demonstrations. ORION first extracts a sequence of Open-World Object Graphs (OOGs), where each OOG models a keyframe state with task-relevant objects and hand information. Then ORION leverages the OOG sequence to construct a manipulation policy that generalizes across varied initial conditions, specifically in four aspects: visual background, camera shifts, spatial layouts, and novel instances from the same object categories.
@@
We introduce our method ORION, short for Open-woRld video ImitatiON. Figure 1 visualizes a high-level overview of ORION. The core innovation lies in creating an object-centric spatiotemporal abstraction that effectively bridges the observational gap between human demonstration and robot execution. The design of ORION stems from our insight that manipulation tasks center around object interaction, and task completion depends on whether specific intermediate states, so-called subgoals, are reached. To capture the object-centric information in the video, we design a graph-based, object-centric representation, called Open-world Object Graphs (OOGs), to model the states of task-relevant objects and their relationships. An OOG has a two-level hierarchy. The high level consists of the object nodes and a hand node, where object nodes identify and localize the relevant objects by leveraging outputs from vision foundation models, while the hand node encodes the interaction information between the hand and objects, such as where to grasp. The low level consists of point nodes, which correspond to object keypoints, and the node features detail the motions of object keypoints in the 3D space.
@@
ORION extracts a manipulation plan from the video as a sequence of OOGs and uses the plan to construct a generalizable policy. Experiments indicate that ORION constructs manipulation policies that are robust to conditions vastly different from the one in the video. Using only an iPhone or an iPad recording of a human performing the task in everyday environments (e.g., an office or a kitchen), the resulting policies succeed in workspaces with drastically different visual backgrounds, camera angles, and spatial arrangements, and even generalize to manipulating unseen object instances of the same categories.
@@
In summary, our contribution is three-fold: 1) We pose the problem of learning vision-based robot manipulation from a single human video in the open-world setting; 2) We introduce Open-world Object Graphs (OOGs), a graph-based, object-centric representation for modeling the states and relations of task-relevant objects; and 3) We present ORION, an algorithm that uses a single video to construct a manipulation policy, which generalizes to conditions that differ in four key ways: visual backgrounds, camera perspectives, spatial configurations, and new object instances.
@@ 2 Problem Formulation
In this paper, we consider a vision-based, tabletop manipulation task, formulated as a finite-horizon Markov Decision Process (MDP) described by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},H,R,\mu\rangle$, where $\mathcal{S}$ is the state space of raw sensory data including RGB-D images and robot proprioception, $\mathcal{A}$ is the action space of low-level robot commands, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ is the transition dynamics, $H$ is the maximal task horizon, $R$ is the sparse reward function, and $\mu$ is the initial state distributions of a task. In this work, we consider the case where task reward functions are defined based on the contact relations between a small set of task-relevant objects. For example, a mug is placed on top of a coaster, or a spoon is put inside a bowl. A reward function returns $1$ if all object relations of a task are satisfied and 00 otherwise. The primary objective of solving a manipulation task is to find a visuomotor policy $\pi$ that maximizes the expected task success rate from a wide range of initial configurations, characterized by $\mu$, where the states vary across the following four dimensions: 1) changing visual backgrounds, 2) different camera angles, 3) different object instances from the same categories, and 4) varied spatial layouts of the task-relevant objects.
@@
We assume a robot does not have direct access to the ground-truth task reward or the physical states of task-relevant objects. We consider a setting where a single actionless video [15, 16] $V$ is provided as a state-only demonstration. We assume $V$ to be a video stream of a person manipulating the task-relevant objects with their single hand, captured as a sequence of RGB-D images using a stationary camera. $V$ is an arbitrarily long video that involves a manipulation sequence where the contact relations among task-relevant objects and the hand change (e.g., an object is grasped or an object is placed on top of another). To avoid the inherent ambiguities of videos due to the distraction of irrelevant objects, each $V$ is accompanied by a complete list of English descriptions of the task-relevant objects, uniquely defining the object instances in $V$. Such a list is represented as a comma-separated list; an example is “[‘small red block’, ‘boat body’]” for the task shown in Figure 2. In this scenario, however, the robot is not pre-programmed to have access to ground-truth categories and locations of the task-relevant objects in $V$. We refer to this challenging setting as “open-world” [17], as the robot must imitate from $V$ while not pre-programmed or trained to interact with the objects in $V$. To allow a robot to operate in this “open-world” setting, we assume access to common sense knowledge through large models pre-trained on internet-scale data, i.e., foundation models.
@@
For evaluation, we adopt the following procedure. Given a single video $V$ that accomplishes a task instance drawn from $\mu$, the performance of an approach is quantified by the average rewards received when evaluating new task instances drawn from the same $\mu$.
@@ 3 Method
In this section, we describe our method ORION (Open-woRld video ImitatiON). ORION is an algorithm that allows a robot to mimic how to perform a manipulation task given a single human video, $V$. To effectively construct a policy $\pi$ from $V$, ORION employs a learning objective based on an object-centric prior. The goal is to create a policy $\pi$ that directs the robot to move objects along 3D trajectories that mimic the directional and curvature patterns observed in $V$, relative to the objects’ initial and final positions. This objective is based on the observation that objects are likely to achieve target configurations by moving along trajectories similar to those in $V$. Key to ORION is generating a manipulation plan from $V$, which serves as the spatiotemporal abstraction of the video that guides the robot to perform a task. A plan is a sequence of object-centric keyframes that each specifies an initial or a subgoal state captured in $V$. We first introduce our formulation of the object-centric representation of a state, Open-world Object Graph (OOG), used in ORION, and then describe the algorithm that constructs a robot policy given a human video.
Figure 2: Overview of plan generation in ORION. ORION generates a manipulation plan from a given video $V$ in order for subsequent policies to synthesize actions. ORION first tracks the objects and keypoints across the video frames. Then keyframes are identified based on the velocity statistics of the keypoint trajectories. Then ORION generates an Open-world Object Graph (OOG) for every keyframe, resulting in a sequence of OOGs that serves as the spatiotemporal abstraction of the video. The figure is viewed best in color.
@@ 3.1 Open-world Object Graph
At the core of our approach is a graph-based, object-centric representation, Open-world Object Graphs (OOGs). OOGs use open-world vision models that model the visual scenes with task-relevant objects and the hand such that they naturally exclude the distracting factors in visual data and localize the task-relevant objects regardless of their spatial locations (see Section 3.2).
@@
We denote an OOG as $\mathcal{G}$. At the high level, each object node corresponds to a task-relevant object from the result of open-world vision models. Every object node comes with node features, consisting of colored 3D point clouds derived from RGB-D observations. This node feature indicates both what and where objects are and also represents their geometry information. Additionally, to inform the robot where to interact with objects (e.g. where to grasp), we introduce the specialized “hand node”, which stores the interaction cues such as contact points and the grip status (open or closed) that can be directly mapped to the robot end-effector during execution. At the low level, each point node corresponds to a keypoint that belongs to a task-relevant object. Every point node comes with the feature, namely the 3D motion trajectories. The feature explicitly models how an object should be moved during a manipulation task. In the rest of the paper, by motion features of a point node in $G_{l}$, we mean 3D trajectory between keyframe $l$ and $l+1$.
@@
In an OOG, all the object nodes and the hand node are fully connected, reflecting real-world spatial relationships. Additionally, the edges are augmented with a binary attribute, indicating whether two objects or objects and the hand are in contact to represent their pairwise contact relations. This attribute allows our developed algorithm to check the set of contact relations that are satisfied, retrieving the matched OOG from the generated plan (see Section 3.2). As for the low-level point nodes, they are connected to their respective object node, indicating a belonging relationship. In the rest of the paper, we denote node entities from human videos with a superscript $V$, and denote the ones from the robot rollout with a superscript $Ro$. Table 1 in the appendix also summarizes the variables needed to define an OOG.
@@ 3.2 Manipulation Plan Generation From $V$
We describe the first part of ORION (see Figure 2), where our method automatically annotates the video and generates a manipulation plan from a given human video, $V$. In this paper, a manipulation plan is a spatiotemporal abstraction of $V$ that centers around the object states and their motions over time, demonstrating how a task should be completed. Our core insight is that we can cost-effectively model a task with object locations at some keyframe states where the set of satisfied contact relations is changed, and abstract a majority of intermediate states into 3D motions of objects. Concretely, a plan is represented as a sequence of OOGs, $\{\mathcal{G}_{l}\}_{l=0}^{L}$ which corresponds to $L+1$ keyframes in $V$, with $\mathcal{G}_{0}$ representing the initial state.
@@
Tracking task-relevant objects. ORION first localizes task-relevant objects in the video $V$. Given $V$ and the list of object descriptions mentioned in Section 2, ORION uses an open-world vision model, Grounded-SAM [18], to annotate video frames with segmentation masks of the task-relevant objects. In practice, open-world vision models are computationally demanding, so we reduce the computation by exploiting object permanence to track the objects. Specifically, ORION annotates the first video frame with Grounded-SAM, and then propagates the segmentation to the rest of the video using a Video Object Segmentation Model, Cutie [19].
@@
Discovering keyframes. After annotating the locations of task-relevant objects, we track their motions across the video to discover the keyframes based on the velocity statistics of object motions. This design is based on the observation that changes in object contact relations due to manipulation are often accompanied by sudden changes in object motions (e.g., transitioning from free space motion to grasping an object). However, keeping full track of object point motion using techniques like optical flow estimation requires heavy computation and the tracking quality is susceptible to noisy observations, largely due to occlusions during manipulation. We use a Track-Any-Point (TAP) model, namely CoTracker [20], to track a subset of points in a long-term video with explicit occlusion modeling, which has been successfully applied to track object motions in robot manipulation [21, 22]. Specifically, we first sample keypoints within the object segmentation of the first frame and track the trajectories across the video. The changes in velocity statistics are straightforward to detect based on the TAP trajectories, where we discover the keyframes using a standard unsupervised changepoint detection algorithm [23].
@@
Generating OOGs from $V$. Once ORION discovers the keyframes, it generates an OOG at each keyframe to model the state of task-relevant objects and the human hand in $V$. The creation of OOG nodes can reuse the results from the annotation process: for object nodes, the point clouds for node features are obtained by back-projecting the object segmentation with depth data; for the point nodes, each node corresponds to the sampled keypoints, and their motion features, 3D trajectories, are back-projected from the TAP trajectories using depth data. Additionally, hand information is required to specify the interaction points with task-relevant objects and the grip status to be mapped to the robot gripper. We use a hand-reconstruction model, HaMeR [24], which gives a reconstructed hand mesh that pinpoints the hand locations at each keyframe. The distances between the fingertips of the mesh help determine the grip status, i.e., whether it is open or closed.
@@
With all the node information, ORION establishes the edge connections between nodes in OOGs, representing contact relations. Since all object and hand locations are computed in the camera frame while the camera extrinsic of $V$ is unknown, there is ambiguity when deciding the spatial relations between objects. We exploit the assumption of tabletop manipulation, where a table is always present with its normal direction aligned with the z-axis of the world coordinate system. So ORION estimates the transformation matrix of the table plane and transforms all the point cloud features in OOGs to align with the xy plane of the world coordinate (Full details appear in Appendix A.2). Then, the contact relations in each state can be determined based on the spatial relations and the computed distances between point clouds. The relations allow ORION to match the test-time observations with a keyframe state from the plan and subsequently decide which object to manipulate (see Section 3.3). In the end, ORION generates a complete OOG for each discovered keyframe.
@@
Figure 3: Overview of the ORION Policy. ORION first localizes task-relevant objects at test time and retrieves the matched OOG from the generated manipulation plan. Then ORION uses the retrieved OOGs to predict the object motions by first computing global registration of object point clouds and then transforming the observed keypoint trajectories from video into the workspace. The predicted trajectories are then used to optimize the SE(3) action sequence of the robot end effector, which is subsequently used to command the robot.

### v2

@@ Abstract
This work presents an object-centric approach to learning vision-based manipulation skills from human videos. We investigate the problem of robot manipulation via imitation in the open-world setting, where a robot learns to manipulate novel objects from a single video demonstration. We introduce ORION, an algorithm that tackles the problem by extracting an object-centric manipulation plan from a single RGB or RGB-D video and deriving a policy that conditions on the extracted plan. Our method enables the robot to learn from videos captured by daily mobile devices and to generalize the policies to deployment environments with varying visual backgrounds, camera angles, spatial layouts, and novel object instances. We systematically evaluate our method on both short-horizon and long-horizon tasks, using RGB-D and RGB-only demonstration videos. Across varied tasks and demonstration types (RGB-D / RGB), we observe an average success rate of 74.4%, demonstrating the efficacy of ORION in learning from a single human video in the open world. Additional materials can be found on the project website.
{NoHyper}*Equal Contribution.
@@ 1 Introduction
A critical step toward building robot autonomy is developing sensorimotor skills for perceiving and interacting with unstructured environments. Conventional methods for acquiring skills necessitate manual engineering and/or costly data collection [1, 2, 3, 4, 5]. A promising alternative is teaching robots through human videos of manipulation behaviors situated in everyday scenarios. These methods have great potential to tap into the readily available source of Internet videos that encompass a wide distribution of human activities, paving the way for scaling up skill learning.
@@
Prior work on learning from human videos has focused on pre-training representations and value functions [6, 7, 8, 9, 10]. However, they do not explicitly capture object interactions and states in 3D space where robot motions are defined. Consequently, they require separate teleoperation data for each set of objects in each location and even for each possible change in visual backgrounds, e.g., the scene background or lighting conditions [11]. In contrast, our goal is for a robot to learn to perform a task robustly in the “open world”, i.e., under varying visual and spatial conditions from a single human video, without prior knowledge of the object models or the behaviors shown. Since our policy construction process uses actionless videos that are equivalent to state-only demonstrations in the problem of “imitation from observation”[12], we refer to our problem setting as open-world imitation from observation.
@@
Developing such a method in this setting becomes possible due to the recent advances in vision foundation models [13, 14]. These models, pre-trained on Internet-scale visual data, excel at understanding open-vocabulary visual concepts and enable robots to recognize and localize objects in natural videos without known object categories or access to physical states. This work marks a major step toward achieving our vision of open-world imitation from observation, where a robot learns to interact with objects given a single video while generalizing to environments with different visual backgrounds and unseen spatial configurations during deployment. In this work, we consider using RGB or RGB-D video demonstrations where a person manipulates a small set of task-relevant objects, recorded with a stationary camera. These videos are actionless or state-only, as they do not come with any ground-truth action labels for the robot.
@@
Figure 1: Overview. We introduce ORION for tackling the problem of learning manipulation behaviors from single human video demonstrations. ORION first extracts a sequence of Open-World Object Graphs (OOGs), where each OOG models a keyframe state with task-relevant objects and hand information. Then, ORION leverages the OOG sequence to construct a manipulation policy that generalizes across varied initial conditions, specifically in four aspects: visual background, camera shifts, spatial layouts, and novel instances from the same object categories.
@@
We introduce our method ORION, short for Open-woRld video ImitatiON. Figure 1 visualizes a high-level overview of ORION. The core innovation lies in creating an object-centric spatiotemporal abstraction that effectively bridges the observational gap between human demonstration and robot execution. The design of ORION stems from our insight that manipulation tasks center around object interaction, and task completion depends on whether specific intermediate states, so-called subgoals, are reached. To capture the object-centric information in the video, we design a graph-based, object-centric representation, called Open-world Object Graphs (OOGs), to model the states of task-relevant objects and their relationships. An OOG has a two-level hierarchy. The high level consists of object nodes and a grasp node, where object nodes identify and localize the relevant objects by leveraging outputs from vision foundation models, while the grasp node encodes the interaction information between the hand and objects, such as where to grasp. The low level consists of point nodes, which correspond to object keypoints, and the node features detail the motions of object keypoints in the 3D space.
@@
ORION extracts a manipulation plan from the video as a sequence of OOGs and uses the plan to construct a generalizable policy. Experiments indicate that ORION constructs manipulation policies that are robust to conditions vastly different from those in the video. Using only an iPhone or iPad recording — or even synthetic videos generated by models such as Veo 2 [15] — of a task being performed in everyday environments (e.g., an office or a kitchen), the resulting policies succeed in workspaces with drastically different visual backgrounds, camera angles, and spatial arrangements, and even generalize to manipulating unseen object instances of the same categories.
@@
In summary, our contribution is threefold. 1) We formulate the problem of learning vision-based robot manipulation from a single human video in the open-world setting; 2) We introduce Open-world Object Graphs (OOGs), a graph-based, object-centric representation that models the states and relations of task-relevant objects; and 3) We present ORION, an algorithm that uses a single video to construct a manipulation policy, which generalizes to conditions that differ in four key ways: visual backgrounds, camera perspectives, spatial configurations, and new object instances.
@@ 2 Problem Formulation
In this paper, we consider a vision-based, tabletop manipulation task, formulated as a finite-horizon Markov Decision Process (MDP) described by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},H,R,\mu\rangle$, where $\mathcal{S}$ is the state space of raw sensory data and robot proprioception, $\mathcal{A}$ is the action space of low-level robot commands, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ is the transition dynamics, $H$ is the maximal task horizon, $R$ is the sparse reward function, and $\mu$ is the initial state distributions of a task. We focus on the case where task reward functions are defined based on the object relations (contact and non-contact relations) between a small set of task-relevant objects. For example, a contact relation change occurs when a mug is placed on top of a coaster. A non-contact relation change occurs when a cup is angled to pour liquid into a bowl. A reward function returns 1 if all object relations of a task are satisfied and 0 otherwise. The primary objective of solving a manipulation task is to find a visuomotor policy $\pi$ that maximizes the expected task success rate from a wide range of initial environment configurations characterized by $\mu$, where the states vary across the following four dimensions: 1) changing visual backgrounds, 2) different camera angles, 3) different object instances from the same categories, and 4) varied spatial layouts of the task-relevant objects.
@@
We assume a robot does not have direct access to the ground-truth task reward or the physical states of task-relevant objects. In addition, we assume a single actionless video [16, 17]. $V$ is provided as a state-only demonstration consisting of an RGB or RGB-D video stream of a person manipulating task-relevant objects, captured as an arbitrarily long sequence of images using a stationary camera. In addition, $V$ involves a manipulation sequence where the object relations among task-relevant objects or hands change (e.g., an object is gripped, then placed on top of another object, or an object is tilted into a pouring position over another object). In this scenario, however, the robot is not pre-programmed to have access to ground-truth categories and locations of the task-relevant objects in $V$. We refer to this challenging setting as “open-world” [18], as the robot must imitate from $V$ while not pre-programmed or trained to interact with the objects in $V$. To allow a robot to operate in this ”open-world” setting, we assume access to common sense knowledge through large models pre-trained on internet-scale data, i.e., foundation models.
@@
For evaluation, we adopt the following procedure. Given a single video $V$ that matches the object relations of a task instance drawn from $\mu$, the model’s performance is measured by the average rewards received when evaluating new task instances drawn from the same $\mu$. If the manipulation in $V$ does not result in object relations that satisfy the task’s requirements (e.g., the task specifies pouring juice into a cup, but the video pours juice on the floor), it is not considered a valid demonstration and thus should not be used for evaluation.
@@ 3 Method
In this section, we describe our method ORION (Open-woRld video ImitatiON). ORION is an algorithm that allows a robot to mimic human execution of a task, given a single demonstration video $V$. To effectively construct a policy $\pi$ from $V$, ORION employs a learning objective based on an object-centric prior. The goal is to create a policy $\pi$ that directs the robot to move objects along trajectories that mimic the directional and curvature patterns observed in $V$, relative to the objects’ initial and final positions. This objective is based on the observation that objects are likely to achieve target configurations by moving along trajectories similar to those in $V$. Key to ORION is generating a manipulation plan from $V$, which serves as the spatiotemporal abstraction of the video that guides the robot to perform a task. A plan is an arbitrarily long sequence of object-centric keyframes, where each keyframe specifies an initial or a subgoal state captured in $V$. We first introduce our formulation of the object-centric representation of a state, Open-world Object Graph (OOG), used in ORION, and then describe the algorithm that constructs a robot policy given a human video.
Figure 2: Overview of plan generation in ORION. ORION generates a manipulation plan from a given video $V$ in order for subsequent policies to synthesize actions. ORION first tracks objects and keypoints across the video frames. Then, keyframes are identified based on the velocity statistics of the keypoint trajectories. Finally, ORION generates an Open-world Object Graph (OOG) for every keyframe, resulting in a sequence of OOGs that serves as the spatiotemporal abstraction of the video. The figure is best viewed in color.
@@ 3.1 Open-world Object Graph
At the core of our approach is a graph-based, object-centric representation, Open-world Object Graphs (OOGs). OOGs use open-world vision models to model visual scenes with task-relevant objects and manipulators (e.g., a human hand or gripper) such that they naturally exclude the distracting factors in visual data and localize the task-relevant objects regardless of their spatial locations (see Section 3.2).
@@
We denote an OOG as $\mathcal{G}$, which consists of object, gripper, and point nodes. At the high level, each object node corresponds to a task-relevant object described by the output of open-world vision models. Every object node comes with node features, consisting of colored 3D point clouds derived from RGB and depth observations. These node features indicate both what and where objects are and also represent their geometry information. For RGB-only videos, an object node additionally includes an estimated object mesh, along with depth estimated from monocular depth estimation models. Next, to inform the robot where to interact with objects (e.g., where to grasp), we introduce the specialized “grasp node”, which stores the interaction cues such as contact points and the grip status (open or closed) that can be directly mapped to the robot end-effector during execution. At the low level, each point node corresponds to a keypoint that belongs to a task-relevant object. Every point node comes with an object-centric feature trajectory. This trajectory contains a list of features describing the object’s state throughout the video, such as a list of 3D points or 6-DOF poses. The information in this trajectory explicitly models how an object should be moved during a manipulation task. In the rest of the paper, by motion features of a point node in $G_{l}$, we mean the object-centric feature trajectory between keyframe $l$ and $l+1$.
@@
In an OOG, all the object nodes and the hand node are fully connected, reflecting real-world spatial relationships. Additionally, the edges are augmented with a binary attribute, indicating whether two objects or an object and the hand are in contact. This attribute allows our developed algorithm to check the set of contact relations that are satisfied, retrieving the matched OOG from the generated plan (see Section 3.2). As for the low-level point nodes, they are connected to their respective object node, indicating a belonging relationship. In the rest of the paper, we denote node entities from human videos with a superscript $V$, and denote the ones from the robot rollout with a superscript $Ro$. Table 1 in the appendix also summarizes the variables needed to define an OOG.
Algorithm 1 Manipulation Plan Generation from Video $V$
1:RGB frames $r_{1...N}$ from demonstration video $V$
2:Depth frames $d_{1..N}$ from demonstration video $V$
3:OOG sequence $\{\mathcal{G}_{0},\dots,\mathcal{G}_{L}\}$
4:// 1. Track task‐relevant objects
[…]
8:if $d\neq\varnothing$ then $\triangleright$ Depth available
9:  $\mathit{keypoints}\leftarrow$ sample_points($\textit{masks}_{1}$)
10:  $\textit{trajectories}\leftarrow$ CoTracker.track(keypoints, $r$) $\triangleright$ Track object keypoints in $V$
[…]
12:  $d\leftarrow$ estimate_depth($r$)
13:  $\mathit{meshes}\leftarrow$ scale_from_depth(InstantMesh($\mathit{mask}$, $r$), $d$) $\triangleright$ Generate object meshes
14:  $\mathit{trajectories}\leftarrow$ FoundationPose($r$, $d$, $\mathit{masks}$, $\mathit{meshes}$) $\triangleright$ Track object poses in $V$
15:end if
16:// 2. Discover keyframes
17:keyframes $\leftarrow$ changepoint_detect(trajectories)
18:// 3. Generate OOGs
[…]
20:for each keyframe $l$ in keyframes do
21:  $\mathcal{G}_{l}\leftarrow$ empty_graph( )
22:  add_object_nodes($\mathcal{G}_{l}$, $\textit{masks}_{l}$, $d_{l}$, $\mathit{meshes}$ if $meshes\neq\varnothing$)
23:  add_point_nodes($\mathcal{G}_{l}$, $\mathit{trajectories}_{l}$)
24:  add_grasp_node($\mathcal{G}_{l}$, $d_{l}$, $\mathit{meshes}$ if $meshes\neq\varnothing$)
25:  $G_{l}\leftarrow$ transform($T,G_{l}$) $\triangleright$ Align OOG features with plane
26:end for
27:return $\{\mathcal{G}_{l}\}_{l=0}^{L}$
@@ 3.2 Manipulation Plan Generation From $V$
We describe the first part of ORION (see Figure 2 and Algorithm 1), where our method automatically annotates the video and generates a manipulation plan from a given human video, $V$. In this paper, a manipulation plan is a spatio-temporal abstraction of $V$ that centers around the object states and their motions over time, demonstrating how a task should be completed. Our core insight is that we can cost-effectively model a task with object locations at some keyframe states where the set of satisfied contact relations is changed, and abstract a majority of intermediate states into object motion trajectories (3D translations or 6-DOF pose trajectories). Concretely, a plan is represented as a sequence of OOGs, $\{\mathcal{G}_{l}\}_{l=0}^{L}$, which corresponds to $L+1$ keyframes in $V$, with $\mathcal{G}_{0}$ representing the initial state. Note that ORION processes demonstration videos differently depending on whether depth information is available. We refer to these as the RGB-D case (with RGB and depth) and the RGB case (with RGB only).
@@
Tracking task-relevant objects. ORION first localizes task-relevant objects in the video $V$ (Lines 1-3). To begin, ORION samples frames from $V$ and passes these frames to a VLM such as GPT-4o. The VLM is then prompted to return a list of strings describing each task-relevant object, while also indicating whether an object is being manipulated (details in Appendix E). This list is then used as an input to an open-world vision model, Grounded-SAM [19], to annotate video frames with segmentation masks of the task-relevant objects. In practice, open-world vision models are computationally demanding, so we reduce the computation by exploiting object permanence to track the objects (Lines 4-10). Specifically, ORION annotates the first video frame with Grounded-SAM, and then propagates the segmentation to the rest of the video using a Video Object Segmentation Model, Cutie [20]. In the RGB case, ORION instead tracks an object’s 6-DOF trajectory with a modified version of FoundationPose [21]. The input to FoundationPose includes estimated depth, per-frame object masks, and an estimated object mesh, which is generated by InstantMesh [22] (implementation details in Appendix A.2).
@@
Discovering keyframes. After annotating the locations of task-relevant objects, ORION tracks their motions across the video to discover $K$ keyframes based on the velocity statistics of object motions (Line 12). This design is based on the observation that changes in object contact relations due to manipulation are often accompanied by sudden changes in object motions (e.g., transitioning from free space motion to grasping an object). However, keeping full track of object point motion using techniques like optical flow estimation often results in unsatisfactory tracking quality, largely due to occlusions during manipulation. ORION uses a track-any-point (TAP) model, namely CoTracker [23], to track a subset of points in a long-term video with explicit occlusion modeling, which has been successfully applied to track object motions in robot manipulation [24, 25]. Specifically, ORION first samples keypoints within the object segmentation of the first frame and track the trajectories across the video. The changes in velocity statistics are straightforward to detect based on the TAP trajectories, where we discover the keyframes using a standard unsupervised changepoint detection algorithm [26]. In the RGB case, we find that setting $K=2$, where the first keyframe is the first video frame and the last keyframe is the last video frame, still results in an OOG sequence that sufficiently describes the object motions in $V$.
@@
Generating OOGs from $V$. Once ORION discovers the keyframes, it generates an OOG at each keyframe to model the state of task-relevant objects and the human hand in $V$. The creation of OOG nodes can reuse the results from the annotation process. For object nodes, the point clouds for node features are obtained by back-projecting the object segmentation with depth data (Line 16). In the RGB-D case, each point node corresponds to sampled keypoints. Their motion features, 3D trajectories, are back-projected from the TAP trajectories using depth data (Line 17). In the RGB case, we have a single point node containing a 6-DOF object trajectory generated by FoundationPose.
Additionally, grip information is required to specify the interaction points with task-relevant objects and the grip status to be mapped to the robot gripper (Line 18). In the RGB-D case, we use a hand-reconstruction model, HaMeR [27], which gives a reconstructed hand mesh that pinpoints the hand locations at each keyframe. The distances between the fingertips of the mesh help determine the grip status, i.e., whether it is open or closed. In the RGB case, since many in-the-wild videos do not show a human hand or significantly occlude the hand, ORION uses grasp heuristics extracted from the estimated object mesh or point cloud, combined with the estimated 6-DOF pose.
@@
With all the node information, ORION establishes the edge connections between nodes in OOGs, representing contact relations. Since all object and hand locations are computed in the camera frame while the camera extrinsic of $V$ is unknown, there is ambiguity when deciding the spatial relations between objects. We exploit the assumption of tabletop manipulation, where a plane is always present with its normal direction aligned with the z-axis of the world coordinate system (Lines 13, 19). ORION estimates the transformation matrix of the table plane and transforms all the point cloud features in OOGs to align with the x-y plane of the world coordinate (Full details appear in Appendix A.2). Then, the contact relations in each state can be determined based on the spatial relations and the computed distances between point clouds. The relations allow ORION to match the test-time observations with a keyframe state from the plan and subsequently decide which object to manipulate (see Section 3.3). In the end, ORION generates a complete OOG for each discovered keyframe.
@@
Figure 3: Overview of the ORION Policy. ORION first localizes task-relevant objects at test time and retrieves the matched OOG from the generated manipulation plan. Then, ORION uses the retrieved OOGs to predict the object motions by warping the object-centric feature trajectory from the video to match the test-time observation. The predicted trajectories are then used to optimize the SE(3) action sequence of the robot end effector, which is subsequently used to command the robot.
Algorithm 2 Robot Policy to Synthesize Actions
1:Manipulation plan $\mathcal{G}=\{\mathcal{G}_{0},\dots,\mathcal{G}_{L}\}$, Demonstration video $V$
2:Sequence of robot end‑effector SE(3) actions, executed on robot
3:while not task_completed and not task_failed do
4:// 1. Retrieve matching OOGs from plan
5:  $r^{Ro}$, $d^{Ro}\leftarrow$ get_robot_obs( )
6:  $\mathcal{G}^{Ro}\leftarrow$ generate_OOG($r^{Ro},d^{Ro}$) $\triangleright$ Same vision pipeline as in  3.2
7:  $(\mathcal{G}_{l},\mathcal{G}_{l+1})\leftarrow$ find_matching_OOG($\mathcal{G}^{Ro},\mathcal{G}$)
8:  $\hat{o}_{target}$, $\hat{o}_{ref}$, $\tau^{V}_{target}\leftarrow$ extract_objs_demo($\mathcal{G}_{l}$)
9:  $o_{target}$, ${o}_{ref}\leftarrow$ extract_objs_test($\hat{o}_{target}$, $\hat{o}_{ref}$, $\mathcal{G}_{l}$, $\mathcal{G}_{l+1}$, $\mathcal{G}^{Ro}$)
10:// 2. Predict object motions
11:  $s_{new},e_{new}=\textsc{find\_rollout\_start\_end}(\hat{o}_{target},{o}_{target},\hat{o}_{ref},{o}_{ref})$
12:  $\tau^{Ro}_{target}\leftarrow$ warp_trajectories($\tau^{V}_{target},s_{new},e_{new}$) $\triangleright$ Details in Algorithm 3 and 4
13:// 3. Optimize robot actions
14:  $\{T_{i}\}_{i=0}^{t_{l+1}-t_{l}}$ $\leftarrow$ $\operatorname*{arg\,min}\limits_{T_{0},T_{1},\dots,T_{t_{l+1}-t_{l}}}\sum_{i=0}^{t_{l+1}-t_{l}}\left\lVert\tau^{Ro}_{\text{target}}({i+1})-T_{i}\tau^{Ro}_{\text{target}}({i})\right\rVert_{2}^{2}$
15:  $\mathit{joint\_seq}\leftarrow$ IK($\{T_{i}\}$)
16:  $\mathit{joint\_seq}\leftarrow$ add_grip_info($\mathit{joint\_seq}$, $\mathcal{G}_{l}.vh$)
17:  robot_execution(joint_seq)
18:end while